2x square + y square+z square +4 yz +2 xy-2 xx




1
Expert's answer
2022-01-21T14:22:15-0500

"2x^{2}+y^{2}+z^{2}+4yz+2xy-2xz"


NOTE: WE SHALL RE-REPRESENT THE VARAIBALE x, y and z TO VARIABLES "x_{1},x_{2}\hspace{0.1cm}and\hspace{0.1cm}x_{3}" respectively.


NOW, using "x_{1},x_{2}\hspace{0.1cm}and\hspace{0.1cm}x_{3}"

"\displaystyle Q(x)=2x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+4x_{2}x_{3}+2x_{1}x_{2}-2x_{1}x_{3}"

The symmetric matrix A with the given quadratic form is

"A=" "\begin{bmatrix}
 2 & 1 & -1 \\
 1 & 1 & 2\\
 -1 & 2 & 1 
\end{bmatrix}"

The characteristic equation of A is

"\mid(A-\lambda I)\mid=-\lambda^{3}+4\lambda^{2}+\lambda-12=0\\\implies \lambda=3,\frac{-1+17}{2},\frac{1+17}{2}"


The corresponding eigen values are:

for "\lambda_{1}=3,\hspace{0.2cm}is\hspace{0.2cm}X_{1}=\begin{bmatrix}
 0 \\
 1\\
 1
\end{bmatrix}"


for "\lambda_{2}=\frac{-1+17}{2},\hspace{0.2cm}is\hspace{0.2cm}X_{2}=\begin{bmatrix}
 \frac{-3+17}{2} \\
 -1\\
 1
\end{bmatrix}"


for "\lambda_{3}=1,\hspace{0.2cm}is\hspace{0.2cm}X_{3}=\begin{bmatrix}
 -\frac{3+17}{2} \\
 -1\\
 1
\end{bmatrix}"


Clearly, the eigen vectors "X_{1}=\begin{bmatrix}
 0 \\
 1\\
 1 
\end{bmatrix}" , "X_{2}=\begin{bmatrix}
 \frac{-3+17}{2} \\
 -1\\
 1
\end{bmatrix}" and "X_{3}=\begin{bmatrix}
 -\frac{3+17}{2} \\
 -1\\
 1
\end{bmatrix}" are linearly independent and orthogonal.